Python OpenSSL Manual

Release 0.9

Martin Sjögren

April 25, 2009

martin@strakt.com

Abstract

This module is a rather thin wrapper around (a subset of) the OpenSSL library. With thin wrapper I mean that a lot of the object methods do nothing more than calling a corresponding function in the OpenSSL library.

Contents

1 Introduction 2

2 Building and Installing 2 2.1 Building the Module on a Unix System . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Building the Module on a Windows System . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3 OpenSSL — Python interface to OpenSSL 3

3.1 crypto — Generic cryptographic module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 X509 objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 X509Name objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 X509Req objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 X509Store objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 PKey objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 PKCS7 objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 PKCS12 objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 X509Extension objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 NetscapeSPKI objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

3.2 rand — An interface to the OpenSSL pseudo random number generator . . . . . . . . . . . . 9 3.3 SSL — An interface to the SSL-specific parts of OpenSSL . . . . . . . . . . . . . . . . . . . . 9

Context objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Connection objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

4 Internals 14 4.1 Exceptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.2 Callbacks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.3 Acessing Socket Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

1 Introduction

The reason pyOpenSSL was created is that the SSL support in the socket module in Python 2.1 (the contemporary version of Python when the pyOpenSSL project was begun) was severely limited. Other OpenSSL wrappers for Python at the time were also limited, though in different ways. Unfortunately, Python’s standard library SSL support has remained weak, although other packages (such as M2Crypto^{1}) have made great advances and now equal or exceed pyOpenSSL’s functionality.

The reason pyOpenSSL continues to be maintained is that there is a significant user community around it, as well as a large amount of software which depends on it. It is a great benefit to many people for pyOpenSSL to continue to exist and advance.

2 Building and Installing

These instructions can also be found in the file INSTALL.

I have tested this on Debian Linux systems (woody and sid), Solaris 2.6 and 2.7. Others have successfully compiled it on Windows and NT.

2.1 Building the Module on a Unix System

pyOpenSSL uses distutils, so there really shouldn’t be any problems. To build the library:

      python setup.py build

If your OpenSSL header files aren’t in /usr/include, you may need to supply the -I flag to let the setup script know where to look. The same goes for the libraries of course, use the -L flag. Note that build won’t accept these flags, so you have to run first build_ext and then build! Example:

      python setup.py build_ext -I/usr/local/ssl/include -L/usr/local/ssl/lib
      python setup.py build

Now you should have a directory called OpenSSL that contains e.g. SSL.so and __init__.py somewhere in the build dicrectory, so just:

      python setup.py install

If you, for some arcane reason, don’t want the module to appear in the site-packages directory, use the

--prefix option.

You can, of course, do

      python setup.py --help

to find out more about how to use the script.

1See http://chandlerproject.org/Projects/MeTooCrypto

2.2 Building the Module on a Windows System

Big thanks to Itamar Shtull-Trauring and Oleg Orlov for their help with Windows build instructions. Same as for Unix systems, we have to separate the build_ext and the build.

Building the library:

      setup.py build_ext -I ...\openssl\inc32 -L ...\openssl\out32dll
      setup.py build

Where ...\openssl is of course the location of your OpenSSL installation.

Installation is the same as for Unix systems:

      setup.py install

And similarily, you can do

      setup.py --help

to get more information.

3 OpenSSL — Python interface to OpenSSL

This package provides a high-level interface to the functions in the OpenSSL library. The following modules are defined:

crypto

Generic cryptographic module. Note that if anything is incomplete, this module is!

rand

An interface to the OpenSSL pseudo random number generator.

SSL

An interface to the SSL-specific parts of OpenSSL.

3.1 crypto — Generic cryptographic module

X509Type

A Python type object representing the X509 object type.

X509()

Factory function that creates an X509 object.

X509NameType

A Python type object representing the X509Name object type.

X509Name(x509name)

Factory function that creates a copy of x509name.

X509ReqType

A Python type object representing the X509Req object type.

X509Req()

Factory function that creates an X509Req object.

X509StoreType

A Python type object representing the X509Store object type.

PKeyType

A Python type object representing the PKey object type.

PKey()

Factory function that creates a PKey object.

PKCS7Type

A Python type object representing the PKCS7 object type.

PKCS12Type

A Python type object representing the PKCS12 object type.

X509ExtensionType

A Python type object representing the X509Extension object type.

X509Extension(typename, critical, value)

Factory function that creates a X509Extension object.

NetscapeSPKIType

A Python type object representing the NetscapeSPKI object type.

NetscapeSPKI([enc ])

Factory function that creates a NetscapeSPKI object. If the enc argument is present, it should be a base64-encoded string representing a NetscapeSPKI object, as returned by the b64_encode method.

FILETYPE_PEM
FILETYPE_ASN1

File type constants.

TYPE_RSA
TYPE_DSA

Key type constants.

exception Error Generic exception used in the crypto module.

dump_certificate(type, cert)

Dump the certificate cert into a buffer string encoded with the type type.

dump_certificate_request(type, req)

Dump the certificate request req into a buffer string encoded with the type type.

dump_privatekey(type, pkey[, cipher, passphrase ])

Dump the private key pkey into a buffer string encoded with the type type, optionally (if type is FILETYPE_PEM) encrypting it using cipher and passphrase.

passphrase must be either a string or a callback for providing the pass phrase.

load_certificate(type, buffer)

Load a certificate (X509) from the string buffer encoded with the type type.

load_certificate_request(type, buffer)

Load a certificate request (X509Req) from the string buffer encoded with the type type.

load_privatekey(type, buffer[, passphrase ])

Load a private key (PKey) from the string buffer encoded with the type type (must be one of

FILETYPE_PEM and FILETYPE_ASN1).

passphrase must be either a string or a callback for providing the pass phrase.

load_pkcs7_data(type, buffer)

Load pkcs7 data from the string buffer encoded with the type type.

load_pkcs12(buffer[, passphrase ])

Load pkcs12 data from the string buffer. If the pkcs12 structure is encrypted, a passphrase must be included.

X509 objects

X509 objects have the following methods:

get_issuer()

Return an X509Name object representing the issuer of the certificate.

get_pubkey()

Return a PKey object representing the public key of the certificate.

get_serial_number()

Return the certificate serial number.

get_subject()

Return an X509Name object representing the subject of the certificate.

get_version()

Return the certificate version.

get_notBefore()

Return a string giving the time before which the certificate is not valid. The string is formatted as an ASN1 GENERALIZEDTIME:

                       YYYYMMDDhhmmssZ
                       YYYYMMDDhhmmss+hhmm
                       YYYYMMDDhhmmss-hhmm

If no value exists for this field, None is returned.

get_notAfter()

Return a string giving the time after which the certificate is not valid. The string is formatted as an ASN1 GENERALIZEDTIME:

                       YYYYMMDDhhmmssZ
                       YYYYMMDDhhmmss+hhmm
                       YYYYMMDDhhmmss-hhmm

If no value exists for this field, None is returned.

set_notBefore(when)

Change the time before which the certificate is not valid. when is a string formatted as an ASN1 GENERALIZEDTIME:

                       YYYYMMDDhhmmssZ
                       YYYYMMDDhhmmss+hhmm
                       YYYYMMDDhhmmss-hhmm
set_notAfter(when)

Change the time after which the certificate is not valid. when is a string formatted as an ASN1 GENERALIZEDTIME:

                            YYYYMMDDhhmmssZ
                            YYYYMMDDhhmmss+hhmm
                            YYYYMMDDhhmmss-hhmm
gmtime_adj_notBefore(time)

Adjust the timestamp (in GMT) when the certificate starts being valid.

gmtime_adj_notAfter(time)

Adjust the timestamp (in GMT) when the certificate stops being valid.

has_expired()

Checks the certificate’s time stamp against current time. Returns true if the certificate has expired and false otherwise.

set_issuer(issuer)

Set the issuer of the certificate to issuer.

set_pubkey(pkey)

Set the public key of the certificate to pkey.

set_serial_number(serialno)

Set the serial number of the certificate to serialno.

set_subject(subject)

Set the subject of the certificate to subject.

set_version(version)

Set the certificate version to version.

sign(pkey, digest)

Sign the certificate, using the key pkey and the message digest algorithm identified by the string digest.

subject_name_hash()

Return the hash of the certificate subject.

digest(digest_name)

Return a digest of the certificate, using the digest_name method. digest_name must be a string describing a digest algorithm supported by OpenSSL (by EVP_get_digestbyname, specifically). For

     example, "md5" or "sha1".
add_extensions(extensions)

Add the extensions in the sequence extensions to the certificate.

X509Name objects

X509Name objects have the following methods:

hash()

Return an integer giving the first four bytes of the MD5 digest of the DER representation of the name.

der()

Return a string giving the DER representation of the name.

get_components()

Return a list of two-tuples of strings giving the components of the name.

X509Name objects have the following members:

countryName

The country of the entity. C may be used as an alias for countryName.

stateOrProvinceName

The state or province of the entity. ST may be used as an alias for stateOrProvinceNameů

localityName

The locality of the entity. L may be used as an alias for localityName.

organizationName

The organization name of the entity. O may be used as an alias for organizationName.

organizationalUnitName

The organizational unit of the entity. OU may be used as an alias for organizationalUnitName.

commonName

The common name of the entity. CN may be used as an alias for commonName.

emailAddress

The e-mail address of the entity.

X509Req objects

X509Req objects have the following methods:

get_pubkey()

Return a PKey object representing the public key of the certificate request.

get_subject()

Return an X509Name object representing the subject of the certificate.

set_pubkey(pkey)

Set the public key of the certificate request to pkey.

sign(pkey, digest)

Sign the certificate request, using the key pkey and the message digest algorithm identified by the string digest.

verify(pkey)

Verify a certificate request using the public key pkey.

set_version(version)

Set the version (RFC 2459, 4.1.2.1) of the certificate request to version.

get_version()

Get the version (RFC 2459, 4.1.2.1) of the certificate request.

X509Store objects

The X509Store object has currently just one method:

add_cert(cert)

Add the certificate cert to the certificate store.

PKey objects

The PKey object has the following methods:

bits()

Return the number of bits of the key.

generate_key(type, bits)

Generate a public/private key pair of the type type (one of TYPE_RSA and TYPE_DSA) with the size bits.

type()

Return the type of the key.

PKCS7 objects

PKCS7 objects have the following methods:

type_is_signed()

FIXME

type_is_enveloped()

FIXME

type_is_signedAndEnveloped()

FIXME

type_is_data()

FIXME

get_type_name()

Get the type name of the PKCS7.

PKCS12 objects

PKCS12 objects have the following methods:

get_certificate()

Return certificate portion of the PKCS12 structure.

get_privatekey()

Return private key portion of the PKCS12 structure

get_ca_certificates()

Return CA certificates within the PKCS12 object as a tuple. Returns None if no CA certificates are present.

X509Extension objects

X509Extension objects have several methods:

get_critical()

Return the critical field of the extension object.

get_short_name()

Return the short type name of the extension object.

NetscapeSPKI objects

NetscapeSPKI objects have the following methods:

b64_encode()

Return a base64-encoded string representation of the object.

get_pubkey()

Return the public key of object.

set_pubkey(key)

Set the public key of the object to key. sign(key, digest_name)

Sign the NetscapeSPKI object using the given key and digest_name. digest_name must be a string describing a digest algorithm supported by OpenSSL (by EVP_get_digestbyname, specifically). For

     example, "md5" or "sha1".
verify(key)

Verify the NetscapeSPKI object using the given key.

3.2 rand — An interface to the OpenSSL pseudo random number generator

This module handles the OpenSSL pseudo random number generator (PRNG) and declares the following:

add(string, entropy)

Mix bytes from string into the PRNG state. The entropy argument is (the lower bound of) an estimate of how much randomness is contained in string, measured in bytes. For more information, see e.g. RFC 1750.

egd(path[, bytes ])

Query the Entropy Gathering Daemon^{2} on socket path for bytes bytes of random data and and uses add to seed the PRNG. The default value of bytes is 255.

load_file(path[, bytes ])

Read bytes bytes (or all of it, if bytes is negative) of data from the file path to seed the PRNG. The default value of bytes is -1.

screen()

Add the current contents of the screen to the PRNG state. Availability: Windows.

seed(string)

This is equivalent to calling add with entropy as the length of the string.

status()

Returns true if the PRNG has been seeded with enough data, and false otherwise.

write_file(path)

Write a number of random bytes (currently 1024) to the file path. This file can then be used with load_file to seed the PRNG again.

3.3 SSL — An interface to the SSL-specific parts of OpenSSL

This module handles things specific to SSL. There are two objects defined: Context, Connection.

SSLv2_METHOD
SSLv3_METHOD
SSLv23_METHOD
TLSv1_METHOD

These constants represent the different SSL methods to use when creating a context object.

VERIFY_NONE
VERIFY_PEER
VERIFY_FAIL_IF_NO_PEER_CERT

These constants represent the verification mode used by the Context object’s set_verify method.

FILETYPE_PEM
FILETYPE_ASN1

File type constants used with the use_certificate_file and use_privatekey_file methods of Context objects.

2See http://www.lothar.com/tech/crypto/

OP_SINGLE_DH_USE
OP_EPHEMERAL_RSA
OP_NO_SSLv2
OP_NO_SSLv3
OP_NO_TLSv1

Constants used with set_options of Context objects. OP_SINGLE_DH_USE means to always create a new key when using ephemeral Diffie-Hellman. OP_EPHEMERAL_RSA means to always use ephemeral RSA keys when doing RSA operations. OP_NO_SSLv2, OP_NO_SSLv3 and OP_NO_TLSv1 means to disable those specific protocols. This is interesting if you’re using e.g. SSLv23_METHOD to get an SSLv2-compatible handshake, but don’t want to use SSLv2.

ContextType

A Python type object representing the Context object type.

Context(method)

Factory function that creates a new Context object given an SSL method. The method should be

SSLv2_METHOD, SSLv3_METHOD, SSLv23_METHOD or TLSv1_METHOD.
ConnectionType

A Python type object representing the Connection object type.

Connection(context, socket)

Factory fucnction that creates a new Connection object given an SSL context and a socket ^{3} object.

exception Error This exception is used as a base class for the other SSL-related exceptions, but may also be raised directly.

Whenever this exception is raised directly, it has a list of error messages from the OpenSSL error queue, where each item is a tuple (lib, function, reason). Here lib, function and reason are all strings, describing where and what the problem is. See err(3) for more information.

exception ZeroReturnError

This exception matches the error return code SSL_ERROR_ZERO_RETURN, and is raised when the SSL Connection has been closed. In SSL 3.0 and TLS 1.0, this only occurs if a closure alert has occurred in the protocol, i.e. the connection has been closed cleanly. Note that this does not necessarily mean that the transport layer (e.g. a socket) has been closed.

It may seem a little strange that this is an exception, but it does match an SSL_ERROR code, and is very convenient.

exception WantReadError

The operation did not complete; the same I/O method should be called again later, with the same arguments. Any I/O method can lead to this since new handshakes can occur at any time.

exception WantWriteError
See WantReadError.
exception WantX509LookupError

The operation did not complete because an application callback has asked to be called again. The I/O method should be called again later, with the same arguments. Note: This won’t occur in this version, as there are no such callbacks in this version.

exception SysCallError

The SysCallError occurs when there’s an I/O error and OpenSSL’s error queue does not contain any information. This can mean two things: An error in the transport protocol, or an end of file that violates the protocol. The parameter to the exception is always a pair (errnum, errstr).

3Actually, all that is required is an object that behaves like a socket, you could even use files, even though it’d be tricky to get the handshakes right!

Context objects

Context objects have the following methods:

check_privatekey()

Check if the private key (loaded with use_privatekey[_file]) matches the certificate (loaded with use_certificate[_file]). Returns None if they match, raises Error otherwise.

get_app_data()

Retrieve application data as set by set_app_data.

get_cert_store()

Retrieve the certificate store (a X509Store object) that the context uses. This can be used to add "trusted" certificates without using the. load_verify_locations() method.

get_timeout()

Retrieve session timeout, as set by set_timeout. The default is 300 seconds.

get_verify_depth()

Retrieve the Context object’s verify depth, as set by set_verify_depth.

get_verify_mode()

Retrieve the Context object’s verify mode, as set by set_verify.

load_client_ca(pemfile)

Read a file with PEM-formatted certificates that will be sent to the client when requesting a client certificate.

load_verify_locations(pemfile, capath)

Specify where CA certificates for verification purposes are located. These are trusted certificates. Note that the certificates have to be in PEM format. If capath is passed, it must be a directory prepared using the c_rehash tool included with OpenSSL. Either, but not both, of pemfile or capath may be

None.
set_default_verify_paths()

Specify that the platform provided CA certificates are to be used for verification purposes. This method may not work properly on OS X.

load_tmp_dh(dhfile)

Load parameters for Ephemeral Diffie-Hellman from dhfile.

set_app_data(data)

Associate data with this Context object. data can be retrieved later using the get_app_data method.

set_cipher_list(ciphers)

Set the list of ciphers to be used in this context. See the OpenSSL manual for more information (e.g. ciphers(1))

set_info_callback(callback)

Set the information callback to callback. This function will be called from time to time during SSL handshakes. callback should take three arguments: a Connection object and two integers. The first integer specifies where in the SSL handshake the function was called, and the other the return code from a (possibly failed) internal function call.

set_options(options)

Add SSL options. Options you have set before are not cleared! This method should be used with the OP_* constants.

set_passwd_cb(callback[, userdata ])

Set the passphrase callback to callback. This function will be called when a private key with a passphrase is loaded. callback must accept three positional arguments. First, an integer giving the maximum length of the passphrase it may return. If the returned passphrase is longer than this, it will be truncated. Second, a boolean value which will be true if the user should be prompted for the passphrase twice and the callback should verify that the two values supplied are equal. Third, the value given as the userdata parameter to set_passwd_cb. If an error occurs, callback should return a false value (e.g. an empty string).

set_session_id(name)

Set the context name within which a session can be reused for this Context object. This is needed when doing session resumption, because there is no way for a stored session to know which Context object it is associated with. name may be any binary data.

set_timeout(timeout)

Set the timeout for newly created sessions for this Context object to timeout. timeout must be given in (whole) seconds. The default value is 300 seconds. See the OpenSSL manual for more information (e.g. SSL_CTX_set_timeout(3)).

set_verify(mode, callback)

Set the verification flags for this Context object to mode and specify that callback should be used for verification callbacks. mode should be one of VERIFY_NONE and VERIFY_PEER. If VERIFY_PEER is

used, mode can be OR:ed with VERIFY_FAIL_IF_NO_PEER_CERT and VERIFY_CLIENT_ONCE to further

control the behaviour. callback should take five arguments: A Connection object, an X509 object, and three integer variables, which are in turn potential error number, error depth and return code. callback should return true if verification passes and false otherwise.

set_verify_depth(depth)

Set the maximum depth for the certificate chain verification that shall be allowed for this Context object.

use_certificate(cert)

Use the certificate cert which has to be a X509 object.

add_extra_chain_cert(cert)

Adds the certificate cert, which has to be a X509 object, to the certificate chain presented together with the certificate.

use_certificate_chain_file(file)

Load a certificate chain from file which must be PEM encoded.

use_privatekey(pkey)

Use the private key pkey which has to be a PKey object.

use_certificate_file(file[, format ])

Load the first certificate found in file. The certificate must be in the format specified by format, which

is either FILETYPE_PEM or FILETYPE_ASN1. The default is FILETYPE_PEM.
use_privatekey_file(file[, format ])

Load the first private key found in file. The private key must be in the format specified by format,

which is either FILETYPE_PEM or FILETYPE_ASN1. The default is FILETYPE_PEM.

Connection objects

Connection objects have the following methods:

accept()

Call the accept method of the underlying socket and set up SSL on the returned socket, using the Context object supplied to this Connection object at creation. Returns a pair (conn, address). where conn is the new Connection object created, and address is as returned by the socket’s accept.

bind(address)

Call the bind method of the underlying socket.

close()

Call the close method of the underlying socket. Note: If you want correct SSL closure, you need to call the shutdown method first.

connect(address)

Call the connect method of the underlying socket and set up SSL on the socket, using the Context object supplied to this Connection object at creation.

connect_ex(address)

Call the connect_ex method of the underlying socket and set up SSL on the socket, using the Context object supplied to this Connection object at creation. Note that if the connect_ex method of the socket doesn’t return 0, SSL won’t be initialized.

do_handshake()

Perform an SSL handshake (usually called after renegotiate or one of set_accept_state or set_accept_state). This can raise the same exceptions as send and recv.

fileno()

Retrieve the file descriptor number for the underlying socket.

listen(backlog)

Call the listen method of the underlying socket.

get_app_data()

Retrieve application data as set by set_app_data.

get_cipher_list()

Retrieve the list of ciphers used by the Connection object. WARNING: This API has changed. It used to take an optional parameter and just return a string, but not it returns the entire list in one go.

get_context()

Retrieve the Context object associated with this Connection.

get_peer_certificate()

Retrieve the other side’s certificate (if any)

getpeername()

Call the getpeername method of the underlying socket.

getsockname()

Call the getsockname method of the underlying socket.

getsockopt(level, optname[, buflen ])

Call the getsockopt method of the underlying socket.

pending()

Retrieve the number of bytes that can be safely read from the SSL buffer (not the underlying transport buffer).

recv(bufsize)

Receive data from the Connection. The return value is a string representing the data received. The maximum amount of data to be received at once, is specified by bufsize.

renegotiate()

Renegotiate the SSL session. Call this if you wish to change cipher suites or anything like that.

send(string)

Send the string data to the Connection.

sendall(string)

Send all of the string data to the Connection. This calls send repeatedly until all data is sent. If an

error occurs, it’s impossible to tell how much data has been sent.

set_accept_state()

Set the connection to work in server mode. The handshake will be handled automatically by read/write.

set_app_data(data)

Associate data with this Connection object. data can be retrieved later using the get_app_data method.

set_connect_state()

Set the connection to work in client mode. The handshake will be handled automatically by read/write.

setblocking(flag)

Call the setblocking method of the underlying socket.

setsockopt(level, optname, value)

Call the setsockopt method of the underlying socket.

shutdown()

Send the shutdown message to the Connection. Returns true if the shutdown message exchange is completed and false otherwise (in which case you call recv() or send() when the connection becomes readable/writeable.

get_shutdown()

Get the shutdown state of the Connection. Returns a bitvector of either or both of SENT_SHUTDOWN and RECEIVED_SHUTDOWN .

set_shutdown(state)

Set the shutdown state of the Connection. state is a bitvector of either or both of SENT_SHUTDOWN and RECEIVED_SHUTDOWN .

sock_shutdown(how)

Call the shutdown method of the underlying socket.

state_string()

Retrieve a verbose string detailing the state of the Connection.

want_read()

Checks if more data has to be read from the transport layer to complete an operation.

want_write()

Checks if there is data to write to the transport layer to complete an operation.

4 Internals

We ran into three main problems developing this: Exceptions, callbacks and accessing socket methods. This is what this chapter is about.

4.1 Exceptions

We realized early that most of the exceptions would be raised by the I/O functions of OpenSSL, so it felt natural to mimic OpenSSL’s error code system, translating them into Python exceptions.

This naturally gives us the exceptions SSL.ZeroReturnError, SSL.WantReadError, SSL.WantWriteError,
SSL.WantX509LookupError and SSL.SysCallError.

For more information about this, see section 3.3.

4.2 Callbacks

There are a number of problems with callbacks. First of all, OpenSSL is written as a C library, it’s not meant to have Python callbacks, so a way around that is needed. Another problem is thread support. A lot of the OpenSSL I/O functions can block if the socket is in blocking mode, and then you want other Python threads to be able to do other things. The real trouble is if you’ve released the global CPython interpreter lock to do a potentially blocking operation, and the operation calls a callback. Then we must take the GIL back, since calling Python APIs without holding it is not allowed.

There are two solutions to the first problem, both of which are necessary. The first solution to use is if the C callback allows ”userdata” to be passed to it (an arbitrary pointer normally). This is great! We can set our Python function object as the real userdata and emulate userdata for the Python function in another way. The other solution can be used if an object with an ”app_data” system always is passed to the callback. For example, the SSL object in OpenSSL has app_data functions and in e.g. the verification callbacks, you can retrieve the related SSL object. What we do is to set our wrapper Connection object as app_data for the SSL object, and we can easily find the Python callback.

The other problem is solved using thread local variables. Whenever the GIL is released before calling into an OpenSSL API, the PyThreadState pointer returned by PyEval_SaveState is stored in a global thread local variable (using Python’s own TLS API, PyThread_set_key_value). When it is necessary to re-acquire the GIL, either after the OpenSSL API returns or in a C callback invoked by that OpenSSL API, the value of the thread local variable is retrieved (PyThread_get_key_value) and used to re-acquire the GIL. This allows Python threads to execute while OpenSSL APIs are running and allows use of any particular pyOpenSSL object from any Python thread, since there is no per-thread state associated with any of these objects and since OpenSSL is threadsafe (as long as properly initialized, as pyOpenSSL initializes it).

4.3 Acessing Socket Methods

We quickly saw the benefit of wrapping socket methods in the SSL.Connection class, for an easy transition into using SSL. The problem here is that the socket module lacks a C API, and all the methods are declared static. One approach would be to have OpenSSL as a submodule to the socket module, placing all the code in ‘socketmodule.c’, but this is obviously not a good solution, since you might not want to import tonnes of extra stuff you’re not going to use when importing the socket module. The other approach is to somehow get a pointer to the method to be called, either the C function, or a callable Python object. This is not really a good solution either, since there’s a lot of lookups involved.

The way it works is that you have to supply a “socket-like” transport object to the SSL.Connection. The only requirement of this object is that it has a fileno() method that returns a file descriptor that’s valid at the C level (i.e. you can use the system calls read and write). If you want to use the connect() or accept() methods of the SSL.Connection object, the transport object has to supply such methods too. Apart from them, any method lookups in the SSL.Connection object that fail are passed on to the underlying transport object.

Future changes might be to allow Python-level transport objects, that instead of having fileno() methods, have read() and write() methods, so more advanced features of Python can be used. This would probably entail some sort of OpenSSL “BIOs”, but converting Python strings back and forth is expensive, so this shouldn’t be used unless necessary. Other nice things would be to be able to pass in different transport objects for reading and writing, but then the fileno() method of SSL.Connection becomes virtually useless. Also, should the method resolution be used on the read-transport or the write-transport?